OpenStudy (anonymous):

A 32.45-g sample of a solid is places in a flask. Toluene, in which the solid is insoluble, is added to the flask so that the total volume of solid and liquid together is 53.00 mL. The solid and toluene together weigh 58.36g. The density of toluene at the temperature of the experiment is 0.864 g/mL. What is the density of the solid?

OpenStudy (anonymous):

Here's what I've got. The toluene must have a mass of 58.36-32.45=25.91 g, right?

OpenStudy (aaronq):

yes, you can then use the density to find the volume it occupies.

OpenStudy (aaronq):

Subtract that you find from 53 mL.

OpenStudy (anonymous):

The volume of the toluene must be grams/density, right?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

So, mL=g/(g/mL)=g*mL/g=mL, yep, mL=25.91/0.864=29.99 mL, right?

OpenStudy (anonymous):

The toluene takes up a volume of 29.99 mL, so the remainder, 53.00-29.99=23.01, must be the solid.

OpenStudy (anonymous):

The solid's mass, 32.45, divided by its volume, 23.01, is its density.

OpenStudy (anonymous):

32.45/23.01=1.410!

OpenStudy (anonymous):

Is that the answer?

OpenStudy (aaronq):

assuming you didn't make any mistakes in the arithmetic, yes it is the answer

OpenStudy (anonymous):

Thanks! It was!

OpenStudy (aaronq):

no problem !